Please help me solve this:
1)Find the solutions of the equation

16x^4=(x-4)^4
2)Find the value of x and y, where x and y are real numbers
(x-y)+3i=7+yi

Answer by edjones(7569)

16x^4=(x-4)^4
4x^2=(x-4)^2
2x=+-(x-4)
.
2x=x-4
x=-4
.
2x=-x+4
3x=4
x=4/3
.
Ed